OpenStudy (anonymous):

The coordinates of rectangle PQRS are P(1,6),Q(1,2),R(9,2) and S(9,6). Show that the diagonals are equal in length.

OpenStudy (johnweldon1993):

Hi :)

OpenStudy (anonymous):

HI!(:

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

Diagonal SQ 

OpenStudy (anonymous):

So are they not of equal lengths?

OpenStudy (johnweldon1993):

They are indeed! ....you'll see :)

OpenStudy (johnweldon1993):

When I put that diagonal...it looks like we have a 2 separate right triangles right?

OpenStudy (anonymous):

Right... So?

OpenStudy (johnweldon1993):

So thanks to the pythagorean theorem...we know that the hypotenuse (diagonal) is equal to \[\large SQ = \sqrt{QR^2 + RS^2} \]

OpenStudy (johnweldon1993):

So what is the distance from point Q to point R ?

OpenStudy (anonymous):

I'm so confused..

OpenStudy (anonymous):

me too but hi :DD

OpenStudy (anonymous):

urhm. Hi.

OpenStudy (jdoe0001):

check how long is from P to R and from Q to S that is, check if their DISTANCE in between those points, the points for the diagonals are the same \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ P&({\color{red}{ 1}}\quad ,&{\color{blue}{ 6}})\quad R&({\color{red}{ 9}}\quad ,&{\color{blue}{ 2}}) \\ \quad \\ Q&({\color{red}{ 1}}\quad ,&{\color{blue}{ 2}})\quad S&({\color{red}{ 9}}\quad ,&{\color{blue}{ 6}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

Wait, so you do the distance formula on PR then QS and thhheeeeennnnn... Compare them?